I could use some help solving this problem please:
Angle H and J are complementary. Angle H measures x+5 and angle J measures 2x-5. Find the measure of angle H.
Please and thank you!

Answer by rfer(16322)

x+5+2x-5=90
3x=90
x=30
x+5=35
2x-5=55